OpenStudy (gibbs):

MEDAL! 1. Nancy has five lbs of mints wich she wants to make 1/4 lb pacages for party flavors. How many pacages can she make? 2. Farmer jones plows 16 2/3 of an acre with his tractor in 20 hours. How long does it take him to plow 1 acre 3. Harry bought a board that was 8ft. long How many 1 1/3 ft. board can cut from the 8 ft. board

OpenStudy (studygurl14):

For the first one: \(\large 5\div\frac{1}{4}\rightarrow 5\times4\rightarrow the~answer\)

OpenStudy (studygurl14):

The second one: \(\large 16\frac{2}{3}\div 20\rightarrow \frac{40}{3}\times\frac{1}{20}\rightarrow\frac{40}{60}\rightarrow the~answer\) (simplify)

OpenStudy (gibbs):

tk and the third one?

OpenStudy (gibbs):

whats the answers

OpenStudy (gibbs):

@StudyGurl14

OpenStudy (studygurl14):

The third one: \(\large 8\div 1\frac{1}{3}\rightarrow 8\times\frac{4}{3}\rightarrow\frac{32}{3}\rightarrow the~answer\) (simplify)